OpenStudy (anonymous):

words problem confuis me a lot plz explain me how to do it plzzz Scott has $500 in the bank. He withdraws the same amount of money each week for 12 weeks, at which point he has $20 left in the bank. How much money did he withdraw each week? A. $40 B. $100 C. $5 D. $60

OpenStudy (anonymous):

i think is a

rishavraj (rishavraj):

@rocky1718 yup./.....

OpenStudy (texaschic101):

500 - 12x = 20 -12x = 20 - 500 -12x = - 480 x = -480/-12 x = 40

OpenStudy (anonymous):

mmm ok

OpenStudy (misty1212):

ooh watch out here @iGreen will report you for giving an answer

OpenStudy (texaschic101):

I did not give an answer...I confirmed an answer

OpenStudy (chosenmatt):

@rocky1718 hi i don't wanna be a bother but don't give direct answers you can get banned http://openstudy.com/code-of-conduct

OpenStudy (igreen):

Exactly. Great work @texaschic101 :D

OpenStudy (chosenmatt):

please don't be mad im just wanna let you know so you can continue using openstudy :) @rocky1718

OpenStudy (texaschic101):

I guess I should have just replied, " you think right ".

OpenStudy (anonymous):

i just put what i think it was i just guees in my own question @chosenmatt

OpenStudy (chosenmatt):

oops i didnt realize it was your question

OpenStudy (chosenmatt):

im soo tired today

OpenStudy (misty1212):

you should ask something ridiculous like "what do you think' as if you were a psychiatrist

OpenStudy (anonymous):

i dont give the answers i give them the promblem after working it out and finding the answer then i gide them though it @misty1212 @iGreen

OpenStudy (texaschic101):

and it just started...you got a long day ahead of you chosenmatt...lol

OpenStudy (chosenmatt):

sorry about that @rocky1718

OpenStudy (anonymous):

is ok @chosenmatt

OpenStudy (chosenmatt):

need help with anything else? @rocky1718 just lemme know :)

OpenStudy (igreen):

That's good @creegan keep it up :)

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

thanks @iGreen

OpenStudy (anonymous):

A plant 3 centimeters tall is placed by a window. If it grows 4 centimeters a day, how many days will it take to be 51 centimeters tall? A. 12 B. 17 C. 12.75 D. 13 @chosenmatt @iGreen

OpenStudy (chosenmatt):

im so embarrassed i said don't give direct answer and it was her own question.....

OpenStudy (chosenmatt):

but anyways one sec

OpenStudy (anonymous):

is ok sweety @chosenmatt

OpenStudy (igreen):

One question per post, close this one and open another one :)

OpenStudy (chosenmatt):

yeah it can cause lag with long threads

OpenStudy (anonymous):

ok